OpenStudy (anonymous):

Graph a hyperbola with foci at (0, ±10) and vertices at (0, ±8) using graphing technology.

OpenStudy (hba):

@brookester6 can help you on this :P

OpenStudy (anonymous):

umm yeah JAS!

OpenStudy (anonymous):

wow a lot of people on one question lol party

OpenStudy (anonymous):

well, I did tag a lot of people. I really need help.

OpenStudy (anonymous):

:( everyone gone.

OpenStudy (loser66):

I witnessed many people stopped by to help you, why did they go all?

OpenStudy (anonymous):

they, left, It's too hard of a question

OpenStudy (anonymous):

@brookester6 said just a second, and then left.

OpenStudy (loser66):

not that girl!! I think because you called too many people. so, this one thinks that the other's duty. hahahaa

OpenStudy (anonymous):

:( aw man, i just need an equation

OpenStudy (anonymous):

Let's start from scratch again

OpenStudy (loser66):

ok, my question is " what does graphing technology?" I always graph by hand, don't know how to use "technology"

OpenStudy (anonymous):

lol using a website, but I just need an equation, I already have a website to use

OpenStudy (loser66):

so, you can solve it by yourself? no need to stay ? right? actually, I am a super lazy guy. If you don't need help anymore, I glad to hear that. hehehe

OpenStudy (anonymous):

no, i need to have the numbers put into an equation, I don't know how to do that

OpenStudy (loser66):

shoot.what is your equation? for hyperbola?

OpenStudy (anonymous):

foci at (0, ±10) and vertices at (0, ±8) Make this into an equation

OpenStudy (loser66):

remember that foci (0,+-c) and vertices (0,+-a) and c^2 = a^2 + b^2( use this form to find out b) and put everything into the standard form of hyperbola \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (loser66):

and draw the asymptotes by \[y= (+-)\frac{a}{b}*x\]

OpenStudy (loser66):

that's it

OpenStudy (anonymous):

10^2=8^2+b^2

OpenStudy (loser66):

remember that if your foci is (0,+- c) your hyperbola has the shape like